OpenStudy (anonymous):

Substitution. x=3 2y+x=3

OpenStudy (texaschic101):

x = 3 now sub 3 in for x in the other equation 2y + x = 3 2y + 3 = 3 --- subtract 3 from both sides 2y = 0 y = 0 check.. 2y + x = 3 2(0) + 3 = 3 3 = 3 (correct) answer : x = 3 and y = 0 any questions ?

OpenStudy (anonymous):

Thank you @texaschic101

OpenStudy (texaschic101):

no problem :)